### v1

# Simul-LLM: A Framework for Exploring High-Quality Simultaneous Translation with Large Language Models

Abstract
Large language models (LLMs) with billions of parameters and pretrained on massive amounts of data are now capable of near or better than state-of-the-art performance in a variety of downstream natural language processing tasks. Neural machine translation (NMT) is one such task that LLMs have been applied to with great success. However, little research has focused on applying LLMs to the more difficult subset of NMT called simultaneous translation (SimulMT), where translation begins before the entire source context is available to the model. In this paper, we address key challenges facing LLMs fine-tuned for SimulMT, validate classical SimulMT concepts and practices in the context of LLMs, explore adapting LLMs that are fine-tuned for NMT to the task of SimulMT, and introduce Simul-LLM111https://github.com/OSU-STARLAB/Simul-LLM, the first open-source fine-tuning and evaluation pipeline development framework for LLMs focused on SimulMT.
Simul-LLM: A Framework for Exploring High-Quality Simultaneous
Translation with Large Language Models
Victor Agostinelli Max Wild Matthew Raffel Kazi Ahmed Asif Fuad Lizhong Chen Oregon State University {agostinv, wildma, raffelm, fuadk, chenliz}@oregonstate.edu
1 Introduction
Modern large language models (LLMs) contain at least several billion and up to trillions of parameters and are remarkably capable across a wide range of tasks. Pretrained on humongous amounts of unlabeled data, they have demonstrated incredible emergent capabilities. With minor prompt adjustments, such as including instructions and examples, LLMs are often capable of near state-of-the-art performance set by highly customized solutions. The performance of these models is further enhanced when fine-tuned for specialized downstream tasks, often exceeding the performance of previously cutting-edge solutions. Given their rapidly evolving capabilities, LLMs and their application have become a focused topic of research within NLP academia.
One popular downstream task for LLMs is text-to-text neural machine translation (NMT), which focuses on taking an input sequence in a given language and outputting a translation in another language. Typically, the entire source context is available at the start of translation for NMT. A particularly challenging subset of NMT is known as simultaneous translation (SimulMT), where the model begins translation without having access to the entire source sequence, and the translation progresses as the remaining source sequence is incrementally provided. For languages that are syntactically and structurally similar, near-NMT performance is fairly achievable, but for language pairs that differ significantly in structure, traditional models struggle to balance high-quality translations with delay for additional source context. This balance is typically achieved via a fixed or adaptive read-write schedule, with one of the most popular and longstanding fixed schedules being the wait-k policy (Ma et al., 2019), where the target translation hypothesis lags behind the incrementally available source sequence by k words or subwords.
While LLMs have been applied to and studied actively in NMT, their application to simultaneous translation has been lagging. This is in part due to a few challenges LLMs face when applied to SimulMT that are non-trivial to address. First and foremost, it is unclear how well LLMs, which are pretrained and usually fine-tuned under the assumption that the prompt is completely provided and static before generation, will adapt to an application space where the prompt dynamically changes as the simultaneous scheduler elects to read from the source sequence. Second, multiple approaches exist to enable LLMs for SimulMT and it is challenging to intuit which approach will perform best. For example, one could adapt LLMs fine-tuned for NMT (hereafter referred to as NMT LLMs) to SimulMT during inference, although how well such models will deal with the source context availability mismatch between fine-tuning (full sentence) and inference (partial sentence) is nebulous. Alternatively, one could fine-tune LLMs directly for SimulMT (hereafter referred to as SimulMT LLMs), but new prompt structuring is likely needed to match inference SimulMT behavior during fine-tuning exactly. Finally, it is also unclear how well previously understood concepts in existing SimulMT work, such as higher fine-tuning wait-k values increasing generalizability, will apply to SimulMT LLMs.
This paper seeks to address the above problems and contributes to the process of applying LLMs to SimulMT in the following major ways:
We develop Simul-LLM, the first open-source fine-tuning and evaluation pipeline development framework for SimulMT LLMs, which seamlessly wraps around and interfaces with popular libraries for LLMs and SimulMT. This framework serves as a foundation for research on SimulMT LLMs that the community can employ and extend for a wide range of future work on LLM-based simultaneous translation.
With the aforementioned framework, we explore the feasibility of adapting LLMs fine-tuned for NMT to SimulMT under a few decoding strategies and the classical wait-k fixed translation scheduler. Generally, we find that NMT LLMs demonstrate good performance during SimulMT inference which can be somewhat boosted by more complex decoding strategies.
We propose an alternative prompt structuring approach to commonly employed NMT prompts that bridges the gap between the fine-tuning and inference environment, assuming a wait-k schedule, and we validate this via the Simul-LLM framework. We elaborate on counter-intuitive results that we observe and provide a base of exploration for future research to employ. Along these lines, we also validate that higher wait-k values employed during SimulMT fine-tuning do increase wait-k generalizability and boost translation quality across the board during SimulMT inference.
2 Background
While a range of work is relevant to this paper, we will only provide a focused and high level review of large language models, LLMs applied towards machine translation, and simultaneous translation as an application space. Readers interested in additional details should engage further with cited works in these areas.
2.1 Large Language Models
Language modeling via transformers, autoregressive (Irie et al., 2019) or bi-directional (Devlin et al., 2019), has been popular nearly since the architecture’s inception, with bi-directional language modeling via BERT demonstrating particularly potent results for its time. In recent years, however, research on other approaches to this task has largely slowed down in favor of rapidly scaling model parameters and employing massive amounts of high-quality data to train on (Kaplan et al., 2020), resulting in the advent of purely generative Large Language Models (LLMs) like GPT-3 (Brown et al., 2020), LLaMa (Touvron et al., 2023), and others.
At their core, LLMs generate the most likely subsequent token (usually at the sub-word level) given all previous tokens in a sequence. Regardless of this simple, core concept, their size and complexity allows for potent pattern recognition and demonstrates numerous emergent capabilities (e.g. simple math, summarization). Additionally, when fine-tuned for specific tasks, this class of models is capable of near-equivalent performance to specialized solutions while, in many cases, maintaining their generalizability to other tasks. For example, this can allow a single end-to-end model to serve as both a high-quality multilingual translation agent and a chat assistant, opening up interesting, practical deployment opportunities.
2.2 Large Language Models for Neural Machine Translation
While LLMs are capable of effectively zero-shot sentence-to-sentence neural machine translation (NMT) (Vilar et al., 2023), their performance can still be improved via simple techniques. Prompt construction has been demonstrated to be critical to LLM performance, both before and after fine-tuning Zhang et al. (2023). One-shot or few-shot performance via In-Context Learning (ICL) can produce near competitive results with fine-tuned LLMs for translation and can even be employed to enhance fine-tuned model performance (Vilar et al., 2023; Xu et al., 2023).
One particularly interesting area of study related to LLMs applied towards NMT remains whether or not to fully fine-tune a given model or engage in Parameter-Efficient Fine-Tuning (PEFT) (Mangrulkar et al., 2022). Early work in this area demonstrated some potential for smaller models (Üstün and Cooper Stickland, 2022), and the accessibility that PEFT provides designers in terms of fine-tuning on low-to-mid performance hardware setups renders it desirable. One of the most popular forms of PEFT freezes a LLM’s weights and adds Low-Rank Adaptation (LoRA) (Hu et al., 2022) adapters between layers (other forms of PEFT exist, but adapter-based PEFT is extremely common so we refer to adapter-based PEFT simply as PEFT hereafter). While fully fine-tuned NMT LLMs tend to suffer from some level of catastrophic-forgetting (Kirkpatrick et al., 2017), intuitively, PEFT-based NMT LLMs should not suffer from any loss of off-task performance, as adapters can be loaded or detached depending on whether or not a given user is prompting for a translation. Given all of these factors, PEFT is an attractive option for NMT LLMs.
2.3 Simultaneous Translation
As a subset of typical NMT, simultaneous translation (SimulMT) focuses on engaging in translation (write decisions) while balancing the amount of available source context (read decisions) to reduce translation latency. This necessarily increases the difficulty of translating a sequence in one language to a sequence in another language, especially when structural and/or syntactical differences exist in the language pair. As a brief example, we can consider translating from a subject-verb-object (SVO) language, such as English, to a subject-object-verb (SOV) language like German. Based on the available source context in English, we may have to guess at the translation in German without access to the necessary context to effectively make that prediction. This is exemplified in Figure 1, where the verb "repaired" must be predicted multiple time-steps before the necessary context is available in the incremental source.
There are two classical, high level approaches to scheduling the write and read decisions of SimulMT, those being static schedules like wait-k (Ma et al., 2019) or adaptive schedules which are flexible and learned, such as variants of monotonic multi-head attention, adaptive wait-k compositions, wait-if-worse, decision state assisted SMT, and others (Grissom II et al., 2014; Gu et al., 2017; Arivazhagan et al., 2019; Zheng et al., 2020). Wait-k remains a particularly popular baseline strategy given its ease of application during training and during inference. It functions by retaining a k-lagging factor between the source context (either in tokens or in words) x and the translation hypothesis y. We can model a typical wait-k schedule’s probability of generating a given output sequence, provided some source sequence, with Equation 1:
Under circumstances or in environments where additional computational latency is acceptable, variations of beam search have been applied in simultaneous scenarios. Speculative Beam Search (SBS) (Zheng et al., 2019) is one such example where, at a high level, each translation step attempts to speculatively translate future steps for some number of beams, eventually selecting a single token or word (the first one) from the most likely beam for some beam length. When applied to wait-k, single token or word SBS can be modeled via Equation 2, where is the length of the beam and represents the speculative beam of maximum joint probability:
Chunk-wise variations are also possible, where for multiple consecutive translation steps, or write decisions, words or subwords from the last determined most-likely beam are employed to cut down on latency.
3 Simul-LLM: an Open-Source SimulMT LLM Fine-tuning Framework
SimulMT is an underexplored application space for LLMs, at the moment, and there is plenty of room to improve performance. To facilitate the rapid development of solutions for SimulMT LLMs (fine-tuned for SimulMT) or NMT LLMs (fine-tuned for NMT) adapted for SimulMT, we choose to develop and provide an open-source framework written in PyTorch for researchers to actively employ for future experiments. We call this framework Simul-LLM, and it bridges the gap between the development of fine-tuning agents via popular libraries and proper SimulMT evaluation. Simul-LLM is poised to support both selective classical SimulMT systems as well as fine-tuning and evaluation for a variety of LLM systems such as Falcon (Almazrouei et al., 2023), LLaMa (Touvron et al., 2023) and Mistral-based (Jiang et al., 2023) models. The high-level components of Simul-LLM include a fine-tuning wrapper and a SimulMT evaluation agent for every supported LLM; in case of classical models, only the evaluation agent is supported in-framework.
BitsAndBytes|.3.1 Fine-tuning Wrapper and Features
The fine-tuning wrapper of Simul-LLM is constructed with a focus on simplicity and extensibility. It is depicted in Figure 2 and supports the following set of user-friendly features.
LLM Support and Extensibility: The proposed Simul-LLM currently supports Falcon and will imminently support Mistral and LLaMa for SimulMT fine-tuning. The fine-tuning wrapper is constructed with a focus on extensibility, allowing for rapid expansion to other LLMs assuming users hold a minimal level of LLM-specific knowledge.
Multiple Prompt Structures: Translation quality, as demonstrated by numerous prior works, varies significantly with prompt structure. As such, the fine-tuning wrapper of Simul-LLM supports multiple kinds of prompt structures, all of which we will validate later in this paper. In the interest of supporting adapting NMT LLMs to SimulMT, Simul-LLM supports NMT fine-tuning in addition to supporting prompt structures that allow for strict wait-k fine-tuning structures.
PEFT and Full Model Fine-tuning: While PEFT (Mangrulkar et al., 2022) has recently become popular, full model fine-tuning is still a valuable option. Simul-LLM supports both PEFT and full model fine-tuning, although PEFT is recommended for most low-to-mid resource hardware setups. We provide recommended PEFT configurations (as seen in Figure 2) for all supported models, and integrate interface for new PEFT configurations when extending to new LLMs.
Flexible Quantization:
Effectively fine-tuning LLMs is reliant on careful memory management for low-to-mid hardware setups. Given that, Simul-LLM quantizes LLMs via the BitsAndBytes
library (seen in the Quant Config in Figure 2), which enables flexible fixed-point quantization. For most low-to-mid hardware setups, we recommend quantizing in 4-bit floating point via 4-bit NormalFloat (nf4).
Prompt Loss Filtering:
Extremely basic supervised fine-tuning may inappropriately include portions of the prompt in loss calculations that are then backpropagated. The fine-tuning wrapper of Simul-LLM ensures that the model only learns from data it is intended to generate post-prompt via a DataCollator
object and a specified response template.
Supervised Fine-tuning Agent:
While custom fine-tuning and optimization loops can be useful, existing and popular options are quite capable. Given that, Simul-LLM’s fine-tuning wrapper surrounds the transformers
supervised fine-tuning trainer and interfaces with it directly. Hyperparameters are fed directly to it as well as PEFT configs, quantization configs, and a specified prompt structure. The trainer itself handles saving model (or LoRA adapter) checkpoints, engaging in validation runs, and optimizing as defined by the provided hyperparameters.
3.2 Evaluation Agent and Features
Evaluation agents for Simul-LLM are similarly built for ease of use and extensibility in addition to customizability towards complex translation schedules and decoding strategies while seamlessly interfacing with the preeminent SimulMT evaluation framework, SimulEval (Ma et al., 2020). This is depicted in Figure 3 and includes the following features.
Classical SimulMT Translation Scheduler: In the interest of baseline accessibility, Simul-LLM evaluation agents support wait-k translation schedules for SimulMT, given their ease of application. Further adaptive or otherwise more involved translation schedulers can be quickly constructed and applied, assuming no reliance on fine-tuning.
Support for Multiple Decoding Strategies: Variable latency constraints for possible inference environments demand flexibility in decoding strategies. As such, evaluation agents for Simul-LLM support several decoding strategies, including greedy and naive decoding, subword-based beam search for single-word decoding, and variations on Speculative Beam Search (SBS) (Zheng et al., 2019), including single word-based SBS and chunk-wise SBS.
Efficient Inference via Custom Generation Stopping Criteria: To avoid excessive generation, Simul-LLM evaluation agents employ custom generation stopping criteria where applicable. This is most relevant in the case of greedy decoding based on singular words, where generation can halt on the first white space delimiter being detected, but constructing more complex criteria is simple given provided examples.
Scoring and Latency via SimulEval: SimulEval (Ma et al., 2020) is the premier SimulMT evaluation framework. Simul-LLM evaluation agents interface seamlessly with SimulEval, which handles the incremental source context and manages translation scoring and latency tracking regarding the translation hypothesis.
4 Adapting NMT LLMs to SimulMT
Existing LLMs that have been fine-tuned for classical NMT may have the potential to be employed directly for SimulMT inference. This can be desirable under circumstances where a single deployed model is preferable to multiple, specialized models (e.g. avoiding fine-tuning costs for multiple models). However, exactly how to adapt such models is unclear in practice given the differences between prompts during NMT fine-tuning (full-sentence availability) and SimulMT inference (incremental source availability). This is especially problematic when engaging with short wait-k schemes and similar low-lagging schedules, where minimal source context is available during early translation steps.
To intuit why this is an issue, suppose that an NMT LLM is accustomed to receiving the entire source sequence x before outputting a word or token . If engaging in a low-lagging wait-k where , then the output can now only be based on source context up to . Assuming the NMT LLM would typically rely on some source context where , then it is missing what should be critical information in its translation decision. We have conducted preliminary quantitative studies on this front by employing the proposed Simul-LLM framework, and these results are presented in Section 7.
5 Prompt Structure for SimulMT LLMs
Alternative to adapting NMT LLMs to SimulMT tasks, the SimulMT LLM approach aims to fine-tune LLMs directly for SimulMT, which requires new prompt structuring. Unlike classical encoder-decoder models for translation, source context for an LLM must be packaged within a prompt and structured appropriately. Some existing work has studied prompt structures for typical NMT (Zhang et al., 2023; Xu et al., 2023), but it is unclear whether such prompts are still optimal for SimulMT. This is because significant differences exist in source context availability between fine-tuning and inference, as discussed in Section 4.
To address this need, we propose a new prompting approach to construct the dataset for fine-tuning and evaluation of SimulMT LLMs. Instead of structuring each source-target sentence pair as a single example as typically done for NMT, we propose decomposing and expanding every sentence pair into a number of examples, where each example incrementally provides additional source and target tokens. When employed, the expanded examples directly mimic the behavior of simultaneous translation. We investigate two specific structures of materializing this prompt structuring approach, as analyzed and compared below.
5.1 Split Source-Target Prompt Structure
The first prompt structure follows the classical NMT prompts where the source sentence is included in the prompt and the target sentence is the output of the LLM. As illustrated in the first prompt structure in Figure 4, with the expanded examples, each example contains partial source along with some instruction in the prompt (starting with "<h>") and partial target that is k words behind the source in the model output (starting with "<a>"), where k is the intended inference wait-k value. While this split source-target prompt structure seems to be a natural and plausible way for model fine-tuning, it has exhibited difficulties when learning for simultaneous translation. The root of this stems from how fine-tuning with LLMs often works: when filtering the prompt from loss calculations during fine-tuning, a response template of some kind (e.g. "<a>:" or ">>ANSWER<<") is employed to ensure only the target word that is generated at the current step (i.e., ) is scored. Unfortunately, with the split source-target prompt, the template allows for all the target words that have been generated from previous translation steps (i.e., to ) to be scored. Without employing a more complex loss filtering, this leads to an inappropriate level of context.
This lack of loss filtering can be especially problematic near the end of a given sequence’s translation. As a simple example, suppose that a given source sequence is of length and . In the first prompt structure where up to source words have been supplied and where , the LLM is effectively being fine-tuned for varying wait-k values ranging from wait-i to wait-k in a single example (as is predicted from to which is wait-k, is predicted from to which is wait-, and so so). If to the point where it is close to normal NMT levels of context, where , then the LLM is no longer being effectively fine-tuned with an appropriate amount of source context for a given translation (write) decision schedule.
5.2 Single Output Word Prompt Structure
The above problem can be entirely side-stepped by our proposed second approach, single output token prompt structure, that embeds only the current target translation hypothesis within the model output.
As illustrated in the second prompt structure in Figure 4, instead of allowing target translation hypotheses from previous time-steps to be incorporated into the loss, the proposed prompt structure shifts those previous translation hypotheses into the prompt. Combined with the expanded examples that form a rigorous wait-k curriculum in terms of the fine-tuning dataset (rigorous meaning a complete curriculum as opposed to a random subset), inference behavior can be copied exactly for every fine-tuning example, thus completely closing the context mismatch between fine-tuning and inference.
Between the two prompt structures, it is clear that the single output word prompt structure more closely replicates the relationship observed in Equation 1 between the source and target sequence during fine-tuning. For both structures, a source-target sentence pair is expanded up to examples.
6 Evaluation Methodology
To validate our proposed solutions to the aforementioned challenges and to test the capabilities of Simul-LLM as the first SimulMT LLM open-source framework, we engaged in several experiments allowing for comparisons among classical non-LLM-based NMT (Vaswani et al., 2017) / SimulMT architectures (Ma et al., 2019), NMT LLMs adapted for SimulMT, and SimulMT LLMs. All mentioned LLMs are fine-tuned Falcon-7B models, but Simul-LLM features an easy to extend framework and we intend to add results for other popular LLMs such as LLaMa-7B and Mistral-7B soon.
Grouped Explorations | Model and Decoding Scheme | en-de | en-es |
Classical Baselines | NMT Transformer (non-simultaneous) | 26.96 (22.6) | 32.64 (23.1) |
Monotonic Transformer Wait-5 (SimulMT) | 22.01 (3.32) | 24.90 (2.58) | |
NMT LLMs Adapted for SimulMT | NMT LLM | 25.83 (3.65) | 30.06 (3.95) |
NMT LLM Single SBS (k=3, b=5, c=1, w=6) | 25.98 (4.12) | 29.48 (4.64) | |
NMT LLM Single SBS (k=3, b=5, c=1, w=10) | 25.95 (4.25) | 27.67 (4.82) | |
NMT LLM Chunk SBS (k=3, b=5, c=2, w=10) | 23.61 (4.63) | 26.33 (5.26) | |
NMT LLM Chunk SBS (k=5, b=5, c=3, w=15) | 25.80 (5.61) | 27.00 (5.90) | |
NMT LLM Chunk SBS (k=7, b=5, c=4, w=20) | 27.32 (6.97) | 28.66 (7.09) | |
SimulMT LLMs with Proposed Prompt | Wait-3 Fine-tuning LLM | 19.99 (3.41) | 23.68 (3.64) |
Wait-7 Fine-tuning LLM | 20.82 (3.44) | 25.18 (3.61) | |
Wait-7 Fine-tuning LLM (k=7) | 23.09 (6.71) | 28.92 (6.87) |
6.1 Dataset Selection and Preprocessing
No standardized dataset exists for SimulMT with LLMs. Due to its popularity in the speech-to-text simultaneous translation (SimulST) track, we employ MuST-C for our experiments222This allows for future work that explores multi-modal simultaneous LLMs, engaging in SimulST via a cascaded model structure with a transcription model for the source speech or a joint speech/text-to-text framework.. For the purposes of adapting MuST-C for text-to-text usage, we preprocess the dataset and filter out certain acoustic indicators (e.g., floating "-" characters representing pauses). In some cases, this resulted in significant changes to some samples of the test set, such as the removal of (Laughter) and (Applause) acoustic indicators that may take up a good portion of the samples.
We employ MuST-C across two language pairs, those being English-to-German (en-de) and English-to-Spanish (en-es). Some additional experiments are provided for the en-es language pair that validate fundamental SimulMT concepts and display BLEU scores, gathered via sacreBLEU (Post, 2018), with respect to samples observed during fine-tuning. The original dataset contains roughly 270K training set samples and approximately 2.5-3K test set samples (tst-COMMON split) per language pair. The expanded version of this dataset for single output word fine-tuning contains approximately 5M training set samples per language pair.
6.2 Training, Fine-Tuning, and Evaluation Hyperparameters and Hardware Details
All classical models were trained on two NVIDIA 32GB V100s and validated on a single V100. All LLMs were fine-tuned via PEFT Mangrulkar et al. (2022) on a single NVIDIA 40GB A40 in bfloat16 and evaluated on a single V100 in float32. Simul-LLM seamlessly integrates with SimulEval (Ma et al., 2020) for the purpose of these evaluations. Classical transformer baselines were trained via Fairseq (Ott et al., 2019), an easily extensible sequence to sequence toolkit.
Classical models were trained with typical hyperparameters provided in Fairseq examples. All LLMs were fine-tuned with identical hyperparameters, employing a constant learning rate of 3e-4 and were optimized via Adam with around 4K warmup updates and batch sizes of 40 samples. LoRA adapter parameters were an of 16 and a value of 64, resulting in a total of around 40M added parameters during fine-tuning, with a dropout value of 0.1. For fair comparison, classical models were of a similar, although not quite identical, size in terms of parameter count.
Additionally, all LLMs were fine-tuned while quantized with NormalFloat4 (nf4) quantization. A small performance boost was observed when removing this quantization during inference, so all models did not engage with nf4 quantization during inference. NMT LLMs were fine-tuned for one epoch, as overfitting was observed beyond that point, which we intuit to be possibly due to the well-documented ability of LLMs to quickly memorize training sets (Biderman et al., 2023). In contrast, SimulMT LLMs were fine-tuned for 2M random examples out of 5M examples on the expanded dataset due to computational constraints.
6.3 Word or Token-Based Wait-k for LLMs
While classical encoder-decoder SimulMT systems usually engage in either word or token-based wait-k, they most typically engage with whichever is more suitable for their vocabulary (i.e. word versus sub-word vocabularies). In spite of the fact that LLMs function via sub-word vocabularies, we recommend, and employ for this work, word-based wait-k for SimulMT LLMs, as it more closely resembles the flow of engaging with a natural language interface. Moreover, supposing that the LLM is receiving a given sequence actively from a transcription system or something similar, it makes intuitive sense to wait for a word to be emitted from the system as opposed to a fragment.
7 Results and Analysis
7.1 Exploration of Adapting NMT LLMs to SimulMT
In Table 1, we provide a breakdown of the performance of several different models, decoding strategies, and wait-k schedules. Regarding our exploration related to adapting NMT LLMs to SimulMT, we also include results related to our implementation of Speculative Beam Search (SBS) (Zheng et al., 2019). As demonstrated by these results, compared with classical models, LLMs fine-tuned for NMT are very capable of SimulMT upon being adapted during inference (even exceeding the score of the classical NMT transformer on en-de that performs non-simultaneous translation). It is worth reiterating that classical architectures typically engage in subword-based wait-k whereas we employ word-based wait-k for LLMs, but the comparisons still serve as a useful reference.
SBS-based decoding strategies helped NMT LLMs in the en-de language pair, but lacked improvement for the en-es language pair. We noted that our implementation (and seemingly also the original implementation) was sensitive to both the window size and the number of committed chunks, with large values for either resulting in the speculative target translation getting too close to the size of the source context. In our tests, when reaching the same length as the source context (akin to context levels of wait-1), degenerate output began to appear that resulted in the output trailing off (e.g., final output of "que…" instead of correct output of "que"). Notably, too many committed chunks only explains performance gaps for chunk-wise SBS, not single SBS, which is normally a flat improvement upon greedy decoding (single SBS is still sensitive to window size). Future experiments can be conducted to utilize the proposed Simul-LLM framework to quantify these factors.
7.2 Exploration of SimulMT LLMs with Proposed Prompt Structure
In Table 1, we also provide a breakdown of the performance of our exploration of SimulMT LLMs with our proposed prompt structure in Section 5.2 that carefully manages source context availability for target translation generation. Two models are employed for this exploration, one fine-tuned for wait-3 inference, and another fine-tuned for wait-7 inference that produces noticeable better quality translations than the first.
While SimulMT LLMs are a more promising approach in achieving higher translation quality than NMT LLMs due to more direct task-specific fine-tuning and better context alignment, our experiment results suggest that, for the time-being, the performance of SimulMT LLMs is not advantageous to NMT LLMs adapted to SimulMT (and varies from slightly worse to barely better translation quality compared with the classical non-LLM SimulMT baseline). This is not completely unexpected as NMT LLMs have been optimized heavily in recent years whereas the exploration of SimulMT LLMs has just started. We provide some analysis below that point out several possible reasons for this observed performance gap and call for additional community efforts to investigate further.
First, due to computational constraints we were unable to fine-tune for an entire epoch of the training dataset (only 2M random samples out of 5M), which represents a major loss of lessons and a lack of rigorous wait-k curriculum for the fine-tuned model. Second, it is possible that the fine-tuning hyperparameters are ill-suited for this particular prompt. We consider this likely to be the most influential issue on our observed results, given the drastic differences between the original and expanded datasets. Third, at least one other work related to NMT LLMs (Chen et al., 2023) has demonstrated that relative positional embeddings can cause issues via attention dilution that ends up being unhelpful, suggesting that distancing the source context, running target hypothesis, and the current translation step hypothesis can be unexpectedly problematic. We posit that our proposed Simul-LLM can be leveraged to verify the above reasons.
Fine-tuning Wait-k | Inference Wait-k | BLEU |
SimulMT LLMs Fine-tuned in Wait-3 | Wait-3 | 23.68 |
Wait-5 | 25.59 | |
Wait-7 | 26.31 | |
SimulMT LLMs Fine-tuned in Wait-7 | Wait-3 | 25.18 |
Wait-5 | 28.19 | |
Wait-7 | 28.92 |
7.3 Higher Wait-k Generalizability Comparisons
It is well documented that in typical SimulMT systems, training or fine-tuning with a slightly higher wait-k than is intended during inference can boost translation quality and generalizability across slightly lower wait-k (Ma et al., 2019). While this likely applies to SimulMT LLMs, no existing work has validated that this behavior persists. We provide a brief comparison of two SimulMT LLMs fine-tuned via wait-3 and wait-7 context levels in Table 2. The results in the table demonstrate that, generally, the expected behavior does hold, with all LLMs fine-tuned in wait-7 outperforming their corresponding wait-3 models for the same inference wait-k. We leave validating additional, previously understood SimulMT principles in SimulMT LLMs to future work.
8 Conclusion
In this work, we introduce Simul-LLM, the first open-source framework that enables rapid development of LLM fine-tuning and evaluation pipelines for simultaneous machine translation (SimulMT). Simul-LLM seamlessly integrates with the fine-tuning and generation tools of the popular transformers
library as well as with SimulEval, the preeminent SimulMT evaluation framework. In addition to introducing Simul-LLM, we employ this framework to explore adapting existing NMT LLMs to SimulMT and propose an expanded fine-tuning dataset and alternative prompt structure to those employed in typical NMT fine-tuning that we believe better replicates inference-time behavior. Moreover, we validate some classically understood SimulMT concepts concerning wait-k scheduling and examine the behavior of SimulMT LLMs during fine-tuning. Our proposed Simul-LLM framework enables multiple lines of future work that can be carried out to understand and optimize LLMs for SimulMT, and it will likely be a useful tool for the research community.